Presented below are certain account balances of Tamarisk Products Co.

Rent revenue

$7,120

Sales discounts

$7,850

Interest expense

13,040

Selling expenses

99,530

Beginning retained earnings

114,740

Sales revenue

405,200

Ending retained earnings

134,890

Income tax expense

24,744

Dividend revenue

71,910

Cost of goods sold

192,416

Sales returns and allowances

12,430

Administrative expenses

74,560
Allocation to noncontrolling interest 19,400


From the foregoing, compute the following: (a) total net revenue, (b) net income, (c) income attributable to controlling stockholders, if Tamarisk has allocation to noncontrolling interest of $19,400.

(a) Total net revenue

$

(b) Net income

$

(c) Income attributable to controlling stockholders

$
Answer #1

a) Total net revenue= sales revenue - Sales discounts - Sales returns and allowances + Rent revenue + Dividend revenue

= 405,200 - 7,850 - 12,430+ 7,120 +71,910

= $463,950

b) Net income= Total net revenue - Cost of goods sold - Administrative expenses - Selling expenses - Interest expense - Income tax expense

= 463,950 - 192,416 - 74,560 - 99,530 - 13,040 - 24,744

= $59,660

c) Income attributable to controlling stockholders = Net income - Allocation to non-controlling interest

= 59,660 - 19,400 = $40,260